OpenStudy (anonymous):

Determine whether the relation shown in the table is a direct variation, an inverse variation, or neither. If it is a direct or inverse variation, write an equation for the function. x ] 10, 15, 20, 25 y ]1/2, 1/3, 1/4, 1/5

OpenStudy (whpalmer4):

Test for indirect variation: x*y = the same value for all values of x and y Test for direct variation: y/x = the same value for all values of x and y

OpenStudy (whpalmer4):

Once you've figured out whether it is direct or indirect variation, you'll either have \[y = kx\]for direct variation, with \(k=\) the same value you found when dividing y/x or \[xy = k\]for indirect variation, with \(k=\) the same value you found when multiplying x*y

OpenStudy (anonymous):

Indirect variation since all the values for x and y were the same. (It did not pass direct variation since the valeus changed.) I'm not sure how to write the equation tho......

OpenStudy (whpalmer4):

Agreed that it is indirect variation, because the product of x and y is always 5. That's how you write the equation: \[xy=5\] Or you can write one of the equivalent forms: \[x=\frac{5}{y}\]\[y=\frac{5}{x}\]

OpenStudy (whpalmer4):

I suppose there might be some situation in which it would be useful to write \[\frac{xy}5=1\]perhaps some situation in which scaling the values was helpful. I haven't ever seen it, but it is equivalent... I would personally go with \(xy=5\) or \(y=\dfrac{5}{x}\)

OpenStudy (anonymous):

I see thank you :)